OpenStudy (anonymous):

solve f(x)+g(x)=6 f(x) = log2(x-7) and g(x) = log2(3x-5) my only answers choices are a) 9 b) -9 c) 512 d) -512

OpenStudy (anonymous):

do this 2 refer to base

OpenStudy (anonymous):

most likely, yes :)

OpenStudy (anonymous):

yes the 2 is the base

OpenStudy (anonymous):

Since the bases of logarithms are matching, we'll apply product rule: log2[(x-7)(3x-5)] = 6 We'll take antilogarithm: [(x-7)(3x-5)] = 2^6 We'll remove brackets: 3x^2 - 26x + 35 = 64 3x^2 - 26x + 35 - 64 = 0

OpenStudy (anonymous):

3x^2 - 26x - 29 = 0 x1 = [26+sqrt(676 +348 )]/6 x1 = (26+32)/6 x1 = 58/6 x1 = 29/3

OpenStudy (anonymous):

x2=-1

OpenStudy (anonymous):

29/3 and -1 are not in the proposed answers...

OpenStudy (anonymous):

there are not correct

OpenStudy (anonymous):

that would explay why i can't find the reight answer to the other question of jaersonandres...

OpenStudy (anonymous):

that would be my teachers fault for putting incorrect answers... thats why i got stuck and asked

OpenStudy (anonymous):

that's crazy that your teacher got wrong in two questions